OpenStudy (anonymous):

x/3 + x/4= 7 FIND X. ~is x 12~

jimthompson5910 (jim_thompson5910):

you are correct, x = 12 is the right answer

jimthompson5910 (jim_thompson5910):

let's check that answer x/3 + x/4= 7 12/3 + 12/4= 7 .. plug in x = 12 4 + 3 = 7 7 = 7 ... that equation is TRUE, so x = 12 is verified to be the true solution

OpenStudy (anonymous):

alright ty

OpenStudy (anonymous):

12/3 = 4 and 12/4 = 3. So 4 + 3 = 7, so yes X would equal 12.

jimthompson5910 (jim_thompson5910):

yw